Shift report 20020116 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Jan 16 2002 - 15:52:04 EST

  • Next message: Flemming Videbaek: "Fw: beam studies"

    ------------------------------------------------------------
    Date: Wed Jan 16 15:52:03 2002, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#12 pages 69 - 70
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam studies until 14:00
       15:00 - Preparing for a RHIC fill
       15:45 - Beam Cogged in RHIC
       15:55 - Beam Dumped (PHENIX & PHOBOS magnet problem ???)
    2. TPC Pedestal   : Run 6612, 6615
       Camac Pedestal : Run 6613
       DVM            : Run 6614, 6617
    
    3. No New detector problems observed.
    
    -------------
     FS moved to "8" degree.
     Magnets D1-D5 set "A"-polarity.
      -> When I asked to change magnet polarity, 
         magnet person found water leak from D2. 
         They replaced hose for D2.
     D1-D4 set 1/8A, and D5 6A.
    
     DAQ looks OK... need to watch "Scale down" for triggers!!!
    
     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Jan 16 2002 - 15:52:46 EST